OpenStudy (anonymous):

What is the sum of the arithmetic sequence 9, 14, 19, …, if there are 34 terms?

OpenStudy (anonymous):

9+(34*5)... maybe

OpenStudy (anonymous):

No, sorry, 9+(33*5)

OpenStudy (anonymous):

Ans Choices A- 3,077 B-3,111 C-3,145 D-3,179

OpenStudy (anonymous):

Yeah nvm what I said earlier

OpenStudy (anonymous):

It would be 9+(5*1)+(5*2)+...(5*33), so you need the sum of 1+2+3+4...+33, and then multiply that by 5, and finally add 9.

OpenStudy (anonymous):

you will get the 34th term by using @Snipa420 's method

OpenStudy (anonymous):

So what is 1+2+3+...33

OpenStudy (anonymous):

Isnt there some math function that makes adding consecutive whole numbers much simpler?

OpenStudy (anonymous):

now use the formula the sum of an a.p = \[n/2(a+t)\]

OpenStudy (anonymous):

where n is the number of terms,a is the 1st term, t is the last term...here t is the 34th term

OpenStudy (anonymous):

Okay so @sritama , 1+2+3+...33=561. 561*5=2805, and 2805+9=2814

OpenStudy (anonymous):

if you solve it,you will get 3111

OpenStudy (anonymous):

thanks (:

OpenStudy (anonymous):

:O...

OpenStudy (anonymous):

@Snipa420 there is no need of adding 1+2+3+....+33 i guess

OpenStudy (anonymous):

I must have done a calculation error.

OpenStudy (anonymous):

What is your method @sritama

OpenStudy (anonymous):

from the formula you wrote first, we will get the 34th term..it is 174

OpenStudy (anonymous):

My second slapjob formula makes sense though...I think.

OpenStudy (anonymous):

so now using the sum of a.p formula, we can write